OpenStudy (vanz):

how do you solve this? 3x^2-24x+48 = 0

OpenStudy (anonymous):

3 (x^2 - 8 + 16) = 0 x^2 -8 + 16 = 0

OpenStudy (anonymous):

(x-4)^2 = 0 x = 4

OpenStudy (vanz):

what happens to the 3? you divide everything by 3?

OpenStudy (anonymous):

0/3 = 0

OpenStudy (vanz):

you divide 3 on both sides?

OpenStudy (anonymous):

yes, ma'am